OpenStudy (anonymous):

City workers are using a snowplow to clear a street. A tire on the snowplow is 3.5ft in diameter and has to turn 36 times in traveling the length of the street. How long is the street? Use the value 3.14 for Pi . Round your answer to the nearest tenth. Do not round any intermediate steps.

OpenStudy (anonymous):

If it turns 36 times to travel the street, that means the street is 36 times the circumference of the tire.

OpenStudy (anonymous):

think you can show me how

OpenStudy (anonymous):

Sure. circumference = pi*diameter circumference = 3.14*3.5 circumference = 10.99. So we know the street is 36 times as long as the circumference of the tire street = 36*circumference street = 36*10.99 street = 395.6ft

OpenStudy (anonymous):

thankyou your a genius can you help me with like two more please